OpenStudy (anonymous):

A chemist has a 30% and a 45% acid solution. What amount of each solution should be used to make 500ml of a solution with 35% acidity?

OpenStudy (anonymous):

\[.45x+.3(500-x)=.35\times 500\] is a good start

OpenStudy (anonymous):

probably easiest to get rid of annoying decimal and solve \[45x+30(500-x)=17500\] btw in this method \(x\) represents the amount of 45% solution